OpenStudy (anonymous):

medal! Directions: Write this expression in expanded form. 3x4y3 + y2 – 2x + 5 A. 3\cdot x\cdot x\cdot x\cdot x\cdot y\cdot y\cdot y\cdot y\cdot y\; +\; 2\; \cdot x\; +\; 5 B. 3\: \cdot \: x\: \cdot \: x\: \cdot \: x\: \cdot y\: \cdot \: y\: \cdot \: y\: +\: y\: \cdot \: y\: -\: 2\: \cdot \: x\: +\: 5 C. 3\: \cdot \: x\: \cdot \: x\: \cdot \: x\: \cdot \: x\: \cdot y\: \cdot y\: \cdot y\: +\: y\: \cdot \: y\: -\: 2\: \cdot \: x\: +\: 5 D. 3\cdot x\cdot\: x\cdot \: x\cdot x\cdot x\cdot y\cdot y\cdot y+\: y\cdot y\; +\; 2\; \cdot x\; +\; 5

OpenStudy (anonymous):

\[A. 3\cdot x\cdot x\cdot x\cdot x\cdot y\cdot y\cdot y\cdot y\cdot y\; +\; 2\; \cdot x\; +\; 5 B. 3\: \cdot \: x\: \cdot \: x\: \cdot \: x\: \cdot y\: \cdot \: y\: \cdot \: y\: +\: y\: \cdot \: y\: -\: 2\: \cdot \: x\: +\: 5 C. 3\: \cdot \: x\: \cdot \: x\: \cdot \: x\: \cdot \: x\: \cdot y\: \cdot y\: \cdot y\: +\: y\: \cdot \: y\: -\: 2\: \cdot \: x\: +\: 5 D. 3\cdot x\cdot\: x\cdot \: x\cdot x\cdot x\cdot y\cdot y\cdot y+\: y\cdot y\; +\; 2\; \cdot x\; +\; 5\]

OpenStudy (anonymous):

wtf is dis

OpenStudy (britchgo):

its 3x^4y^3-2x+y^2+5

OpenStudy (britchgo):

so... 3 x x x x y y y y y-2 x+5

OpenStudy (anonymous):

thanks

OpenStudy (britchgo):

no problem